OpenStudy (anonymous):

Factor. 4x2 + 10x A. 2x2(2x + 5) B. x(2x + 5) C. 4x(x + 5) D. 2x(2x + 5)

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

well I think its A but idk

OpenStudy (skullpatrol):

First of all when you type: 4x2 + 10x, do you mean 4x^2 + 10x which means: $$ 4x^2 + 10x$$

OpenStudy (anonymous):

yeah 4x^2 + 10x

OpenStudy (skullpatrol):

OK, so let's look for "common" factors ...

OpenStudy (skullpatrol):

$$4x^2 + 10x =2*2x^2 + 2*5x $$

OpenStudy (skullpatrol):

What do the two terms on the right hand side have in common?

OpenStudy (anonymous):

I don't really know .. I never got to study this cause my teacher skipped it and now I have a huge test about it ...

OpenStudy (skullpatrol):

What do they share?

OpenStudy (skullpatrol):

The share the 2 and an x, right?

OpenStudy (skullpatrol):

$$\Huge 4x^2 + 10x =2*2x^2+2*5x$$

OpenStudy (skullpatrol):